# An extension of Steinberg’s Theorem to biquotient pairs of subgroups

### v2

@@ 2.3 Enlarging tori while keeping intersections almost intact
Suppose $T_{1}$ and $T_{2}^{(1)},\dots,T_{2}^{(n)}$ are subtori of a torus $T$, with all $T_{2}^{(i)}$ of equal rank. Let $r:=\operatorname{rank}T$, $r_{1}:=\operatorname{rank}K_{1}$, $r_{2}:=\operatorname{rank}K_{2}^{(i)}$, $r^{\prime}:=\max_{i}\operatorname{rank}(T_{1}\cap T_{2}^{(i)})$ and $r_{1}^{+}:=r-r_{2}+r^{\prime}$. There exists a torus $T_{1}^{+}$ such that $T_{1}\subseteq T_{1}^{+}\subseteq T$, $r_{1}^{+}=\operatorname{rank}T_{1}^{+}$ and, for each $i$, $r^{\prime}=\operatorname{rank}(T_{1}^{+}\cap T_{2}^{(i)})$.
This clearly proves the claim we set out to establish. ∎
@@ 2.4 Reduction to toroidal subgroups
In case $H_{1}$ and $H_{2}$ are good, Theorem 13 can be used to show that the implication of the proposition is even an equivalence, but we will not need this.
@@
By symmetry, it suffices to show how to pass from $T_{1}$ to $H_{1}$: we will argue that $\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}T_{1},\mathrm{R}H_{2})=0$ implies $\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}H_{1},\mathrm{R}H_{2})=0$. As in the proof [SCSQ, Corollary 3.10] (the corollary quoted as Corollary 18 above), this can be seen using holomorphic induction (Theorem 10): the inclusion $i\colon T_{1}\hookrightarrow H_{1}$ defines a homomorphism of $\mathrm{R}H_{1}$-modules $i_{*}$ splitting $i^{*}$, and homomorphisms of $\mathrm{R}H_{1}$-modules are, in particular, homomorphisms of $\mathrm{R}G$-modules. So the claim follows. Localization does not affect this argument. ∎
@@ 3 The proof
We now proceed to prove Theorem 3, starting from the most special case and generalizing in several steps. We repeatedly use the fact that $\operatorname{Tor}$ commutes with localization: for any prime ideal $\mathfrak{p}\subset\mathrm{R}G$, we have canonical isomorphisms
$[\operatorname{Tor}_{\mathrm{R}G}^{i}(\mathrm{R}H_{1},\mathrm{R}H_{2})]_{\mathfrak{p}}\cong\operatorname{Tor}_{(\mathrm{R}G)_{\mathfrak{p}}}^{i}((\mathrm{R}H_{1})_{\mathfrak{p}},(\mathrm{R}H_{2})_{\mathfrak{p}}).$ (1)
Recall from Proposition 22 (d) that $\mathrm{Z}G$ denotes the centre of $G$.
@@
Assume that $H_{1}$ is a torus. Assume moreover that the restriction $\mathrm{R}G\to\mathrm{R}H_{2}$ is surjective and that its kernel is generated by a regular sequence $\mathbb{y}=(y_{1},\dots,y_{c_{2}})$ of length $c_{2}=\operatorname{rank}G-\operatorname{rank}H_{2}$. Then $\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}H_{1},\mathrm{R}H_{2})_{\mathfrak{p}}=0$ for all $i>0$ and for all prime ideals $\mathfrak{p}\in\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}(\mathrm{Z}G)$.
@@
Choose a maximal torus $T\subset G$ containing the torus $H_{1}$, and let $i\colon T\hookrightarrow G$ denote the inclusion. Then $H_{1}$ is a direct factor of $T$ [HilgertNeeb]*Lem. 15.3.2, and the kernel of the restriction $\mathrm{R}T\to\mathrm{R}H_{1}$ is generated by a regular sequence $\mathbb{x}=(x_{1},\dots,x_{c_{1}})$ of length $c_{1}=\operatorname{rank}G-\operatorname{rank}H_{1}$. It follows that $\mathrm{R}H_{1}\cong\mathrm{R}T/(\mathbb{x})$ as an $\mathrm{R}T$-module, and that the Koszul complex defined by $\mathbb{x}$ provides a free resolution of $\mathrm{R}H_{1}$ as an $\mathrm{R}T$-module. We write $\Lambda^{\bullet}_{\mathrm{R}T}(\mathrm{R}T;\mathbb{x})$ for this complex, and $\Lambda^{\bullet}_{\mathrm{R}T}(\mathrm{R}T;\mathbb{x})\to\mathrm{R}H_{1}$ for the resolution.
@@
For $H_{2}$, we have an isomorphism of $\mathrm{R}G$-modules $\mathrm{R}H_{2}\cong\mathrm{R}G/(\mathbb{y})$ by assumption, and the Koszul complex $\Lambda^{\bullet}_{\mathrm{R}G}(\mathrm{R}G;\mathbb{y})$ provides a free resolution of $\mathrm{R}H_{2}$ as an $\mathrm{R}G$-module.
@@
Now observe that we have an isomorphism of $\mathrm{R}G$-modules, and an isomorphism of complexes of $\mathrm{R}G$-modules, as follows:
$\mathrm{R}H_{1}\otimes_{\mathrm{R}G}\mathrm{R}H_{2}$ $\cong\mathrm{R}T/(\mathbb{x},i^{*}\mathbb{y})$ (3)
$\Lambda^{\bullet}_{\mathrm{R}T}(\mathrm{R}T;\mathbb{x})\otimes_{\mathrm{R}G}\Lambda^{\bullet}_{\mathrm{R}G}(\mathrm{R}G;\mathbb{y})$ $\cong\Lambda^{\bullet}_{\mathrm{R}T}(\mathrm{R}T;\mathbb{x},i^{*}\mathbb{y})$ (4)
@@
The first isomorphism is clear. To verify the second isomorphism, we can decompose each Koszul complex into a tensor product of Koszul complexes of lengths two:
@@
$\Lambda^{\bullet}_{\mathrm{R}T}(\mathrm{R}T;\mathbb{x})$ $\underset{\mathrm{R}T}{\cong}\bigotimes_{\begin{subarray}{c}\mathrm{R}T\\ i=1,\dots,c_{1}\end{subarray}}(\mathrm{R}T\xrightarrow{x_{i}}\mathrm{R}T)$
@@
$\Lambda^{\bullet}_{\mathrm{R}G}(\mathrm{R}G;\mathbb{y})$ $\underset{\mathrm{R}G}{\cong}\bigotimes_{\begin{subarray}{c}\mathrm{R}G\\ j=1,\dots,c_{2}\end{subarray}}(\mathrm{R}G\xrightarrow{y_{j}}\mathrm{R}G)$
@@
The isomorphism (4) can now be obtained by successively applying the isomorphisms
[…]
@@
We have taken care to distinguish $\mathrm{R}T$-module isomorphisms from $\mathrm{R}G$-module isomorphisms in the notation, but note that every isomorphism of $\mathrm{R}T$-modules is a fortiori an isomorphism of $\mathrm{R}G$-modules.
Viewing the left side of (4) as a double complex, we obtain an associated spectral sequence of $\mathrm{R}G$-modules:
$E_{0}^{p,q}=\Lambda^{p}_{\mathrm{R}T}(\mathrm{R}T;\mathbb{x})\otimes_{\mathrm{R}G}\Lambda^{q}_{\mathrm{R}G}(\mathrm{R}G;\mathbb{y})\quad\Rightarrow\quad\mathrm{H}^{p+q}(\Lambda^{\bullet}_{\mathrm{R}T}(\mathrm{R}T;\mathbb{x},i^{*}\mathbb{y})),$
As $\Lambda^{\bullet}_{\mathrm{R}T}(\mathrm{R}T;\mathbb{x})$ is a resolution of $\mathrm{R}H_{1}$, and as each term of $\Lambda^{\bullet}_{\mathrm{R}G}(\mathrm{R}G;\mathbb{y})$ is a free $\mathrm{R}G$-module, the $E_{1}$-page is concentrated in a single column as follows:
$E^{p,q}_{1}=\begin{cases}\mathrm{R}H_{1}\otimes_{\mathrm{R}G}\Lambda^{q}_{\mathrm{R}G}(\mathrm{R}G;\mathbb{y})&\text{ for }p=0\\ 0&\text{ for }p\neq 0\end{cases}$
As $\Lambda^{\bullet}_{\mathrm{R}G}(\mathrm{R}G;\mathbb{y})$ is a free resolution of $\mathrm{R}H_{2}$, the $E_{2}$-page computes $\operatorname{Tor}$-groups as follows:
$E^{p,q}_{2}=\begin{cases}\operatorname{Tor}_{RG}^{q}(\mathrm{R}H_{1},\mathrm{R}H_{2})&\text{ for }p=0\\ 0&\text{ for }p\neq 0\end{cases}$
At this stage, the spectral sequence necessarily collapses, i. e. $E^{p,q}_{2}=E^{p,q}_{\infty}$. We thus obtain, for all $i$, an isomorphism of $\mathrm{R}G$-modules:
$\operatorname{Tor}_{\mathrm{R}G}^{i}(\mathrm{R}H_{1},\mathrm{R}H_{2})\cong\mathrm{H}^{i}(\Lambda^{\bullet}_{\mathrm{R}T}(\mathrm{R}T;\mathbb{x},i^{*}\mathbb{y}))$ (5)
@@
The basic strategy now is to compute the Krull dimension of $\mathrm{R}T/(\mathbb{x},i^{*}\mathbb{y})$, to deduce that $(\mathbb{x},i^{*}\mathbb{y})$ is a regular sequence in $\mathrm{R}T$, and to conclude that the augmented complex $\Lambda^{\bullet}_{\mathrm{R}T}(\mathrm{R}T;\mathbb{x},i^{*}\mathbb{y})\to\mathrm{R}T/(\mathbb{x},i^{*}\mathbb{y})$ is exact, so that the cohomology groups in (5) vanish in positive degrees.
@@
In order to implement this strategy, we now fix and localize at a prime ideal $\mathfrak{p}$ of $\mathrm{R}G$ that lies in $\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}(\mathrm{Z}G)$. In view of eq. 1, we may and will moreover assume that $\mathfrak{p}\in\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}H_{1}\cap\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}H_{2}$.
@@
By Proposition 22 (d), there is a unique ideal $\mathfrak{p}_{T}\subset\mathrm{R}T$ lying over $\mathfrak{p}\subset\mathrm{R}G$. It follows from the construction that $(\mathbb{x},i^{*}\mathbb{y})\subseteq\mathfrak{p}_{T}$: We have $(\mathbb{x})=\ker(\mathrm{R}T\to\mathrm{R}H_{1})$, so $(i^{*})^{-1}(\mathbb{x})\subseteq\ker(\mathrm{R}G\to\mathrm{R}H_{1})$. Moreover, by the equivalence ($b\Leftrightarrow c$) of Proposition 21, $\ker(\mathrm{R}G\to\mathrm{R}H_{1})\subseteq\mathfrak{p}$. So $(i^{*})^{-1}(\mathbb{x})\subseteq\mathfrak{p}$, and we deduce from Lemma 8 (c) that $(\mathbb{x})\subseteq\mathfrak{p}_{T}$. Similarly, $(\mathbb{y})=\ker(\mathrm{R}G\to\mathrm{R}H_{2})\subseteq\mathfrak{p}$, and as $\mathfrak{p}=(i^{*})^{-1}(\mathfrak{p}_{T})$, we find $(i^{*}\mathbb{y})\subseteq\mathfrak{p}_{T}$.
@@
We can therefore view $(\mathbb{x},i^{*}\mathbb{y})$ as a sequence in the maximal ideal $\mathfrak{p}_{T}\cdot(\mathrm{R}T)_{\mathfrak{p}}$ of the local ring $(\mathrm{R}T)_{\mathfrak{p}}$. Using (3) and the first equality of Proposition 22 (b), we see that $\mathfrak{p}\in\mathrm{Supp}_{\mathrm{R}G}(\mathrm{R}T/(\mathbb{x},i^{*}\mathbb{y}))$. By Lemma 12, we can find another prime $\mathfrak{p}^{\prime}\in\mathrm{Supp}_{\mathrm{R}G}(\mathrm{R}T/(\mathbb{x},i^{*}\mathbb{y}))$ with $\mathfrak{p}^{\prime}\subseteq\mathfrak{p}$ such that the first equality in the following chain of equalities and inequalities holds:
$\dim(\mathrm{R}T)_{\mathfrak{p}}-\dim(\mathrm{R}T)_{\mathfrak{p}}/(\mathbb{x},i^{*}\mathbb{y})$ $=\dim(\mathrm{R}G)_{\mathfrak{p}}-\dim(\mathrm{R}G/\mathfrak{p}^{\prime})_{\mathfrak{p}}$
$=\dim(\mathrm{R}G)-\dim(\mathrm{R}G/\mathfrak{p}^{\prime})$

### v3

@@ 2.3 Enlarging tori while keeping intersections almost intact
(T) Suppose $T_{1}$ and $T_{2}^{(1)},\dots,T_{2}^{(n)}$ are subtori of a torus $T$, with all $T_{2}^{(i)}$ of equal rank. Let $r:=\operatorname{rank}T$, $r_{1}:=\operatorname{rank}T_{1}$, $r_{2}:=\operatorname{rank}T_{2}^{(i)}$, $r^{\prime}:=\max_{i}\operatorname{rank}(T_{1}\cap T_{2}^{(i)})$ and $r_{1}^{+}:=r-r_{2}+r^{\prime}$. There exists a torus $T_{1}^{+}$ such that $T_{1}\subseteq T_{1}^{+}\subseteq T$, $r_{1}^{+}=\operatorname{rank}T_{1}^{+}$ and, for each $i$, $r^{\prime}=\operatorname{rank}(T_{1}^{+}\cap T_{2}^{(i)})$.
Using Lemma˜4, (T) may be translated into the following equivalent statement:
(Z) Suppose $K_{1}$ and $K_{2}^{(1)},\dots,K_{2}^{(n)}$ are direct summands of $\mathbb{Z}^{r}$, with all $K_{2}^{(i)}$ of equal rank. Let $c_{1}:=\operatorname{rank}K_{1}$, $c_{2}:=\operatorname{rank}K_{2}^{(i)}$, $c^{\prime}:=\min_{i}\operatorname{rank}(K_{1}+K_{2}^{(i)})$ and $c_{1}^{-}:=c^{\prime}-c_{2}$. There exists a direct summand $K_{1}^{-}$ of $\mathbb{Z}^{r}$ such that $K_{1}^{-}\subseteq K_{1}$ with $c_{1}^{-}=\operatorname{rank}K_{1}^{-}$ and, for each $i$, $c^{\prime}=\operatorname{rank}(K_{1}^{-}+K_{2}^{(i)})$.
We claim that it suffices to prove the following analogous statement (Q) about rational vector spaces:
(Q) Suppose $W_{1}$ and $W_{2}^{(1)},\dots,W_{2}^{(n)}$ are vector subspaces of $\mathbb{Q}^{r}$, with all $W_{2}^{(i)}$ of equal dimension. Let $c_{1}:=\dim W_{1}$, $c_{2}:=\dim W_{2}^{(i)}$, $c^{\prime}:=\min_{i}\dim(W_{1}+W_{2}^{(i)})$ and $c_{1}^{-}:=c^{\prime}-c_{2}$. There exists a vector subspace $W_{1}^{-}\subseteq W_{1}$ such that $c_{1}^{-}=\dim W_{1}^{-}$ and, for each $i$, $c^{\prime}=\dim(W_{1}^{-}+W_{2}^{(i)})$.
To deduce (Z) from (Q), note that there is a close relation between vector subspaces of $\mathbb{Q}^{r}$ and subgroups of $\mathbb{Z}^{r}$. For a vector subspace $W\subseteq\mathbb{Q}^{r}$, the intersection $W\cap\mathbb{Z}^{r}$ is a direct summand of $\mathbb{Z}^{r}$, of rank equal to the dimension of $W$. Conversely, from any subgroup $K\subseteq\mathbb{Z}^{r}$, we can obtain the vector subspace $K\otimes_{\mathbb{Z}}\mathbb{Q}\subseteq\mathbb{Q}^{r}$ of dimension equal to the rank of $K$. We obtain in this way a bijection between vector subspaces of $\mathbb{Q}^{r}$ and direct summands of $\mathbb{Z}^{r}$. Moreover, for arbitrary subgroups $K_{1},K_{2}\subseteq\mathbb{Z}^{r}$, we have $(K_{1}+K_{2})\otimes\mathbb{Q}=K_{1}\otimes\mathbb{Q}+K_{2}\otimes\mathbb{Q}$. (Note that $K_{1}+K_{2}$ may not be a direct summand even when $K_{1}$ and $K_{2}$ are.) Thus, given direct summands $K_{1}$ and $K_{2}^{(i)}$ as in (Z), we may consider the vector spaces $W_{1}:=K_{1}\otimes\mathbb{Q}$ and $W_{2}^{(i)}:=K_{2}^{(i)}\otimes\mathbb{Q}$ in (Q), obtain a vector space $W_{1}^{-}\subseteq W_{1}$, and then take $K_{1}^{-}:=W_{1}^{-}\cap\mathbb{Z}^{r}$. As noted, this is a direct summand of $\mathbb{Z}^{r}$, clearly contained in $K_{1}$, and the ranks work out as intended.
Next, we reduce (Q) to the case $n=1$:
(Q) Let $W_{1}$ and $W_{2}$ be vector subspaces of $\mathbb{Q}^{r}$, with $c_{i}:=\dim W_{i}$. For any integer $c_{1}^{-}$ such that $0\leq c_{1}^{-}\leq\dim(W_{1}+W_{2})-c_{2}$, there exists a vector subspace $W_{1}^{-}\subseteq W_{1}$ of dimension $c_{1}^{-}$ such that $\dim(W_{1}^{-}+W_{2})=c_{1}^{-}+c_{2}$. Moreover, the vector subspaces $W_{1}^{-}$ with these properties form a (non-empty) Zariski-open subset of the rational Grassmannian $\mathrm{Gr}_{\mathbb{Q}}(c_{1}^{-},W_{1})$.
To deduce (Q) from (Q), note that for each $i$ we have $c_{2}\leq c^{\prime}\leq\dim(W_{1}+W_{2}^{(i)})$ in (Q), so $0\leq c_{1}^{-}\leq\dim(W_{1}+W_{2}^{(i)})-c_{2}$. Thus, by (Q), for each $i$ the vector subspaces $W_{1}^{-}\subseteq W_{1}$ with $\dim(W_{1}^{-}+W_{2}^{(i)})=c^{\prime}$ form a non-empty Zariski-open subset of $\mathrm{Gr}_{\mathbb{Q}}(c_{1}^{-},W_{1})$. So we can pick any $W_{1}^{-}$ that lies in the intersection of these finitely many Zariski open subsets.
Finally, (Q) can be phrased more elegantly, and proved, by noting that $\dim(W_{1}+W_{2})-c_{2}=c_{1}-\dim(W_{1}\cap W_{2})$, and by noting that the condition on the dimension of $W_{1}^{-}+W_{2}$ is equivalent to the condition $W_{1}^{-}\cap W_{2}=\{0\}$. Such a $W_{1}^{-}$ exists because we can simply choose a complement $W^{\prime}$ of $W_{1}\cap W_{2}$ in $W_{1}$, and then pick an arbitrary subspace $W_{1}^{-}\subseteq W^{\prime}$ of dimension $c_{1}^{-}$. To see that trivial intersection with $W_{2}$ is an open condition on $\mathrm{Gr}_{\mathbb{Q}}(c_{1}^{-},W_{1})$, note that this condition on $W_{1}^{-}$ can be rephrased as follows: the restriction of the natural projection $W_{1}\twoheadrightarrow W_{1}/(W_{1}\cap W_{2})$ to $W_{1}^{-}$ is a monomorphism. By change of coordinates, we may assume without loss of generality that this projection is the projection onto the first $\dim(W_{1}+W_{2})-c_{2}$ coordinates. Then the restriction is a monomorphism if and only if, for a given basis $(b_{1},\dots,b_{c_{1}^{-}})$ of $W_{1}^{-}$, one of the maximal minors is non-zero. These $(c_{1}^{-}\times c_{1}^{-})$-minors are precisely the Plücker coordinates on the Grassmannian, so this is an open condition. ∎
@@ 2.4 Reduction to toroidal subgroups
In case $H_{1}$ and $H_{2}$ are good, Theorem˜14 can be used to show that the implication of the proposition is even an equivalence, but we will not need this.
@@
By symmetry, it suffices to show how to pass from $T_{1}$ to $H_{1}$: we will argue that $\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}T_{1},\mathrm{R}H_{2})=0$ implies $\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}H_{1},\mathrm{R}H_{2})=0$. As in the proof of [SCSQ, Corollary 3.10] (the corollary quoted as Corollary˜19 above), this can be seen using holomorphic induction (Theorem˜11): the inclusion $i\colon T_{1}\hookrightarrow H_{1}$ defines a homomorphism of $\mathrm{R}H_{1}$-modules $i_{*}$ splitting $i^{*}$, and homomorphisms of $\mathrm{R}H_{1}$-modules are, in particular, homomorphisms of $\mathrm{R}G$-modules. So $i^{*}$ induces a split injection of $\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}H_{1},\mathrm{R}H_{2})$ into $\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}T_{1},\mathrm{R}H_{2})$, and the claim follows. Localization does not affect this argument. ∎
@@ 3 The proof
We now proceed to prove Theorem˜3, starting from the most special case and generalizing in several steps. We repeatedly use the fact that $\operatorname{Tor}$ commutes with localization: for any prime ideal $\mathfrak{p}\subset\mathrm{R}G$, we have canonical isomorphisms
$\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}H_{1},\mathrm{R}H_{2})]_{\mathfrak{p}}\cong\operatorname{Tor}_{i}^{(\mathrm{R}G)_{\mathfrak{p}}}((\mathrm{R}H_{1})_{\mathfrak{p}},(\mathrm{R}H_{2})_{\mathfrak{p}}).$ (1)
Recall from Proposition˜23 (d) that $\mathrm{Z}G$ denotes the centre of $G$.
@@
Assume that $H_{1}$ is a torus. Assume moreover that the restriction $\mathrm{R}G\to\mathrm{R}H_{2}$ is surjective and that its kernel is generated by a regular sequence $\mathbf{y}=(y_{1},\dots,y_{c_{2}})$ of length $c_{2}=\operatorname{rank}G-\operatorname{rank}H_{2}$. Then $\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}H_{1},\mathrm{R}H_{2})_{\mathfrak{p}}=0$ for all $i>0$ and for all prime ideals $\mathfrak{p}\in\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}(\mathrm{Z}G)$.
@@
Choose a maximal torus $T\subset G$ containing the torus $H_{1}$, and let $i\colon T\hookrightarrow G$ denote the inclusion. Then $H_{1}$ is a direct factor of $T$ [HilgertNeeb]*Lem. 15.3.2, and the kernel of the restriction $\mathrm{R}T\to\mathrm{R}H_{1}$ is generated by a regular sequence $\mathbf{x}=(x_{1},\dots,x_{c_{1}})$ of length $c_{1}=\operatorname{rank}G-\operatorname{rank}H_{1}$. It follows that $\mathrm{R}H_{1}\cong\mathrm{R}T/(\mathbf{x})$ as an $\mathrm{R}T$-module, and that the Koszul complex defined by $\mathbf{x}$ provides a free resolution of $\mathrm{R}H_{1}$ as an $\mathrm{R}T$-module. We write $\Lambda^{\bullet}_{\mathrm{R}T}(\mathbf{x})$ for this complex, and $\Lambda^{\bullet}_{\mathrm{R}T}(\mathbf{x})\to\mathrm{R}H_{1}$ for the resolution. As $\mathrm{R}T$ is a free $\mathrm{R}G$-module, this is also a free resolution of $\mathrm{R}H_{1}$ as an $\mathrm{R}G$-module.
@@
For $H_{2}$, we have an isomorphism of $\mathrm{R}G$-modules $\mathrm{R}H_{2}\cong\mathrm{R}G/(\mathbf{y})$ by assumption, and the Koszul complex $\Lambda^{\bullet}_{\mathrm{R}G}(\mathbf{y})$ provides a free resolution of $\mathrm{R}H_{2}$ as an $\mathrm{R}G$-module.
With these free resolutions $\Lambda^{\bullet}_{\mathrm{R}T}(\mathbf{x})$ and $\Lambda^{\bullet}_{\mathrm{R}G}(\mathbf{y})$ of $\mathrm{R}H_{1}$ and $\mathrm{R}H_{2}$ at our disposal, we can compute the Tor groups we are interested in as the homology groups of the tensor product of the resolutions (see, for example, [CartanEilenberg, Chapter VI, above Proposition 1.2]):
$\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}H_{1},\mathrm{R}H_{2})\cong\mathrm{H}_{i}(\Lambda^{\bullet}_{\mathrm{R}T}(\mathbf{x})\otimes_{\mathrm{R}G}\Lambda^{\bullet}_{\mathrm{R}G}(\mathbf{y}))$ (3)
@@
Next, observe that we have an isomorphism of $\mathrm{R}G$-modules, and an isomorphism of complexes of $\mathrm{R}G$-modules, as follows:
$\mathrm{R}H_{1}\otimes_{\mathrm{R}G}\mathrm{R}H_{2}$ $\cong\mathrm{R}T/(\mathbf{x},i^{*}\mathbf{y})$ (4)
$\Lambda^{\bullet}_{\mathrm{R}T}(\mathbf{x})\otimes_{\mathrm{R}G}\Lambda^{\bullet}_{\mathrm{R}G}(\mathbf{y})$ $\cong\Lambda^{\bullet}_{\mathrm{R}T}(\mathbf{x},i^{*}\mathbf{y})$ (5)
@@
The first isomorphism is clear. To verify the second isomorphism, note that Koszul complexes can be decomposed into tensor products of Koszul complexes of lengths two. Specifically, in our case, we have
@@
$\Lambda^{\bullet}_{\mathrm{R}G}(\mathbf{y})$ $\underset{\mathrm{R}G}{\cong}\;\bigotimes_{\mathclap{\begin{subarray}{c}\mathrm{R}G\\ j=1,\dots,c_{2}\end{subarray}}}\left(\mathrm{R}G\xrightarrow{y_{j}}\mathrm{R}G\right),$
and similarly
@@
$\Lambda^{\bullet}_{\mathrm{R}T}(\mathbf{x},i^{*}\mathbf{y})$ $\underset{\mathrm{R}T}{\cong}\;\Lambda^{\bullet}_{\mathrm{R}T}(\mathbf{x})\;\otimes_{\mathrm{R}T}\;\bigotimes_{\mathclap{\begin{subarray}{c}\mathrm{R}G\\ j=1,\dots,c_{2}\end{subarray}}}\left(\mathrm{R}T\xrightarrow{i^{*}y_{j}}\mathrm{R}T\right).$
@@
In view of these decompositions, the isomorphism (5) can be verified by successively applying the isomorphisms
$\left(\mathrm{R}T\xrightarrow{i^{*}y_{j}}\mathrm{R}T\right)$ $\underset{\mathrm{R}G}{\cong}\mathrm{R}T\otimes_{\mathrm{R}G}(\mathrm{R}G\xrightarrow{y_{j}}\mathrm{R}G)$
@@
for $j=1,\dots,c_{2}$. We have taken care to distinguish $\mathrm{R}T$-module isomorphisms from $\mathrm{R}G$-module isomorphisms in the notation, but note that every isomorphism of $\mathrm{R}T$-modules is a fortiori an isomorphism of $\mathrm{R}G$-modules.
Combining (3) with (5), we arrive at the following isomorphism of $\mathrm{R}G$-modules:
$\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}H_{1},\mathrm{R}H_{2})\cong\mathrm{H}_{i}(\Lambda^{\bullet}_{\mathrm{R}T}(\mathbf{x},i^{*}\mathbf{y})).$ (6)
@@
The basic strategy now is to compute the Krull dimension of $\mathrm{R}T/(\mathbf{x},i^{*}\mathbf{y})$, to deduce that $(\mathbf{x},i^{*}\mathbf{y})$ is a regular sequence in $\mathrm{R}T$, and to conclude that the augmented complex $\Lambda^{\bullet}_{\mathrm{R}T}(\mathbf{x},i^{*}\mathbf{y})\to\mathrm{R}T/(\mathbf{x},i^{*}\mathbf{y})$ is exact, so that the homology groups in (6) vanish in positive degrees.
@@
In order to implement this strategy, we now fix and localize at a prime ideal $\mathfrak{p}$ of $\mathrm{R}G$ that lies in $\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}(\mathrm{Z}G)$. In view of eq.˜1, we may and will moreover assume that $\mathfrak{p}\in\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}H_{1}\cap\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}H_{2}$.
@@
By Proposition˜23 (d), there is a unique ideal $\mathfrak{p}_{T}\subset\mathrm{R}T$ lying over $\mathfrak{p}\subset\mathrm{R}G$. It follows from the construction that $(\mathbf{x},i^{*}\mathbf{y})\subseteq\mathfrak{p}_{T}$: We have $(\mathbf{x})=\ker(\mathrm{R}T\to\mathrm{R}H_{1})$, so $(i^{*})^{-1}(\mathbf{x})\subseteq\ker(\mathrm{R}G\to\mathrm{R}H_{1})$. Moreover, by the equivalence (b $\Leftrightarrow$ c) of Proposition˜22, $\ker(\mathrm{R}G\to\mathrm{R}H_{1})\subseteq\mathfrak{p}$. So $(i^{*})^{-1}(\mathbf{x})\subseteq\mathfrak{p}$, and we deduce from Lemma˜9 (c) that $(\mathbf{x})\subseteq\mathfrak{p}_{T}$. Similarly, $(\mathbf{y})=\ker(\mathrm{R}G\to\mathrm{R}H_{2})\subseteq\mathfrak{p}$, and as $\mathfrak{p}=(i^{*})^{-1}(\mathfrak{p}_{T})$, we find $(i^{*}\mathbf{y})\subseteq\mathfrak{p}_{T}$.
@@
We can therefore view $(\mathbf{x},i^{*}\mathbf{y})$ as a sequence in the maximal ideal $\mathfrak{p}_{T}\cdot(\mathrm{R}T)_{\mathfrak{p}}$ of the local ring $(\mathrm{R}T)_{\mathfrak{p}}$. Using (4) and the first equality of Proposition˜23 (b), we see that $\mathfrak{p}\in\mathrm{Supp}_{\mathrm{R}G}(\mathrm{R}T/(\mathbf{x},i^{*}\mathbf{y}))$. By Lemma˜13, we can find another prime $\mathfrak{p}^{\prime}\in\mathrm{Supp}_{\mathrm{R}G}(\mathrm{R}T/(\mathbf{x},i^{*}\mathbf{y}))$ with $\mathfrak{p}^{\prime}\subseteq\mathfrak{p}$ such that the first equality in the following chain of equalities and inequalities holds:
$\dim{(\mathrm{R}T)_{\mathfrak{p}}}-\dim{(\mathrm{R}T/(\mathbf{x},i^{*}\mathbf{y}))_{\mathfrak{p}}}$ $=\dim{(\mathrm{R}G)_{\mathfrak{p}}}-\dim{(\mathrm{R}G/\mathfrak{p}^{\prime})_{\mathfrak{p}}}$ (7)
$=\dim{\mathrm{R}G}-\dim{\mathrm{R}G/\mathfrak{p}^{\prime}}$